# Evolutionary Tracks and Spectral Properties of Quasi-stars and Their Correlation with Little Red Dots

### v1

@@ Abstract
JWST has revealed a population of red, compact, high-redshift () objects referred to as “Little Red Dots” (LRDs). These objects exhibit unusual spectral features reminiscent of stellar spectra with blackbody-like SEDs, large hydrogen Balmer breaks, Balmer line absorption, and classical stellar absorption features such as calcium H&K and the calcium triplet. Following the recent suggestion that these may be actively accreting direct-collapse black holes in the process of assembly, i.e. quasi-stars, we present evolutionary models of quasi-stars using our recently released, publicly available MESA-QUEST modeling framework. We compute a grid of models spanning a range of black hole masses and predict the luminosities, temperatures, surface gravities, and lifetimes of these objects. We find that these models lie along a Hayashi track once they hit their “late-stage” which constitutes the majority of their lives ( Myr). We present scaling relations for estimating the mass of a quasi-star as a function of the bolometric luminosity, as well as the bolometric luminosity as a function of the effective temperature for the Hayashi track. The short lifetimes in tandem with the observed number density of LRDs imply the possibility that every supermassive black hole was once a quasi-star. We compare synthetic spectra of our quasi-star models to observations of LRDs, and show that these models are broadly capable of reproducing the continuum spectra of observed LRDs. These results indicate that quasi-stars are promising candidates for the origin of supermassive black holes via direct collapse in the early universe.
@@ 1 Introduction
The rapid emergence of supermassive black holes (SMBHs) within the first Myr of cosmic history remains a major puzzle. Observations indicating black hole-to-stellar mass ratios up to , significantly higher than those observed locally (Y. Harikane et al., 2023; B. L. Jones et al., 2025; R. Maiolino et al., 2025; M. Yue et al., 2024), suggest that the theoretically proposed “direct collapse” pathway for forming massive initial black hole seeds may have been in operation in the early Universe (D. J. Whalen et al., 2023; P. Natarajan et al., 2024; E. R. Coughlin & M. C. Begelman, 2024). In this scenario, a heavy seed black hole of order M⊙ forms from the gravitational collapse of low metallicity gas in early-universe dense pre-galactic gas clouds in various cosmic settings, which then accretes at the Eddington limit until it fully forms a SMBH (M. G. Haehnelt & M. J. Rees, 1993; A. Loeb & F. A. Rasio, 1994; D. J. Eisenstein & A. Loeb, 1995; P. Madau & M. J. Rees, 2001; V. Bromm & A. Loeb, 2003; G. Lodato & P. Natarajan, 2006; M. C. Begelman et al., 2006; P. Natarajan et al., 2017) or from the rapid amplification of a light initial seed that can grow at super-Eddington rates (B. Devecchi & M. Volonteri, 2009; L. Mayer et al., 2010; T. Alexander & P. Natarajan, 2014).
@@
The evolution of direct collapse heavy seed formation is expected to proceed in the form of a quasi-star (G. Lodato & P. Natarajan, 2006; M. C. Begelman et al., 2008). These are formed when the core of either a supermassive star or a pre-galactic gas cloud directly collapses initially into a stellar mass black hole around which the remaining gas is assembled as an envelope. This envelope around the growing black hole seed is then supported by the radiation pressure produced from accretion onto the black hole, which can rapidly accrete at super-Eddington rates due to the efficient transport of energy via convection and jets, allowing the central black hole to accrete at roughly the Eddington limit of the entire quasi-star (M. C. Begelman et al., 2006, 2008; M. C. Begelman, 2010; W. H. Ball, 2012a; D. Fiacconi & E. M. Rossi, 2015; C. B. Campbell et al., 2025).
@@
Among the high-redshift, actively-accreting black hole observations by JWST are a subset of candidate active galactic nuclei (AGN) that have come to be known as “Little Red Dots” (LRDs) – high-redshift, morphologically compact, ultra-luminous objects with unique spectral characteristics (J. Matthee et al., 2024; J. E. Greene et al., 2024; I. Labbe et al., 2025; H. B. Akins et al., 2024; I. Juodžbalis et al., 2024; V. Kokorev et al., 2024; D. D. Kocevski et al., 2025; Hviding, Raphael E. et al., 2025; X. Ji et al., 2025; A. J. Taylor et al., 2025). These objects exhibit an unusually strong Balmer break as well as an extended, flat spectrum in the near-infrared (NIR). The combination of these features, in addition to a blue UV slope, results in a unique “V”-shaped spectral energy distribution (SED, G. Barro et al., 2024; Killi, Meghana et al., 2024; de Graaff, Anna et al., 2025; D. D. Kocevski et al., 2025; I. Labbe et al., 2023).
@@
Currently, there is no complete explanation for what these LRDs are composed of and hence their unique signatures. The relative contributions of emitted flux originating from star formation versus accretion onto a central BH are debated. Several studies suggest that LRDs are black holes embedded in dense gas, with R. P. Naidu et al. (2025) dubbing their model a “black hole star” (e.g., de Graaff, Anna et al., 2025; K. Inayoshi & R. Maiolino, 2025; X. Ji et al., 2025; D. Kido et al., 2025; H. Liu et al., 2025; A. J. Taylor et al., 2025). This is motivated by their nearly blackbody-like SEDs, prominent hydrogen Balmer breaks, Balmer line absorption, and classical stellar absorption features such as Ca H&K and the Ca II triplet, which are consistent with emission from an optically thick, thermally emitting envelope surrounding an accreting black hole. J. Jeon et al. (2025) further proposed that LRDs could be remnants of direct collapse, consistent with the quasi-star scenario (M. C. Begelman et al., 2006, 2008; M. C. Begelman, 2010), and M. C. Begelman & J. Dexter (2025) have argued that they would most likely represent the “late stage” of this process where the BH has accreted at least 10% the total mass of the system.
@@
Thus far, there has been no comparison between quasi-star evolutionary models and JWST’s LRDs. In this work, we present simulations of quasi-star evolution using MESA (B. Paxton et al., 2011, 2013, 2015, 2018, 2019; A. S. Jermyn et al., 2023) and the MESA-QUEST modeling framework (C. B. Campbell et al., 2025; A. D. Santarelli et al., 2025). We extract spectral predictions to compare with recent JWST observations of two LRDs: MoM-BH*-1 (R. P. Naidu et al., 2025) and UNCOVER-45924 (I. Labbe et al., 2024). We compare our predictions with observations for both a quasi-star by itself and a quasi-star embedded within a star-forming galaxy, akin to the scenario predicted by P. Natarajan et al. (2017). We also consider simple corrections for dust attenuation, but find little evidence for it, in line with recent findings that dust attenuation in LRDs may be subtle or negligible (C. M. Casey et al., 2025; de Graaff, Anna et al., 2025; R. P. Naidu et al., 2025; D. J. Setton et al., 2025; Xiao, Mengyuan et al., 2025). We find that our fiducial M⊙ quasi-star model broadly reproduces the continuum spectra of these LRDs, thereby lending evidence that SMBHs may assemble by passing through the direct-collapse quasi-star phase.
@@
The outline of our paper is as follows: in Section 2, we provide details of how we extract spectra from quasi-star models in MESA-QUEST. We present the results from models and comparisons with observational data of two JWST LRDs in Section 3, and close with a final discussion and implications of these results.
@@
2 Methods: extracting quasi-star spectra from MESA-QUEST
@@
In this section, we briefly outline our quasi-star MESA models. We use our previously developed framework MESA-QUEST, which includes several newly implemented upgrades: flexible inner boundary conditions, accretion schemes, and improvements that enable modeling higher total masses than previously studied. The following section briefly outlines these methods; more details can be found in A. D. Santarelli et al. (2025) and C. B. Campbell et al. (2025). We note that all MESA models and plotting scripts are publicly available via the MESA-QUEST Github repository.111http://www.github.com/andysantarelli/MESA-QUEST
2.1 Quasi-star Modeling
@@
A quasi-star is modelled as black hole enveloped within a star via an adjustment to the inner boundary conditions of the stellar structure equations such that:
@@
where is the innermost grid point, is the updated inner boundary, and and are the black hole mass and accretion luminosity respectively. In this work, we use the “saturated-convection” radius, the boundary within which material is slowly drifting inward while convection transports energy outward at the highest possible rate (E. R. Coughlin & M. C. Begelman, 2024). Details on the calculation of this boundary can be found in E. R. Coughlin & M. C. Begelman (2024), with further implementation details in A. D. Santarelli et al. (2025) and C. B. Campbell et al. (2025). For the accretion luminosity, and thus the accretion rate, we use the Eddington limit of the entire object such that
@@
Previous works modeling the accretion and resulting luminosity report only a factor of deviation from the object’s Eddington limit throughout the quasi-star’s lifetime (W. H. Ball, 2012b; C. B. Campbell et al., 2025). The dimensionless scaling factor serves as a proxy for a range of physical effects that enhance or suppress accretion such as rotation, magnetic fields, photon-trapping, etc., which we plan to implement and explore more thoroughly in future work. Here we adopt a simple scaling factor and use as a first approximation.
@@
after which the quasi-star “settles” as the black hole mass reaches M⊙. The inset plot shows the majority of the quasi-star’s life as it dims and cools.
The final age of this model is Myr, though the actual lifetimes may be longer as our models do not currently indicate what happens once the inner boundary condition exceeds the total radius. If we assume that BH growth continues at a similar rate until the entire object is consumed, we expect a total lifetime of Myr.
@@
where for convenience we have dropped the dependency on and , though we note that their variation can have a large influence on the accretion physics and thus the luminosity itself (A. D. Santarelli et al., 2025).
2.2 Extracting Synthetic Spectral-Energy Distributions
@@
In order to extract the synthetic spectra of quasi-stars, we use the MESA colors module, developed by N. J. Miller (2025). This is a built-in tool within MESA that can be used to generate synthetic spectral-energy distributions (SEDs) by sampling the nearest grid point in a stellar atmosphere model grid using the stellar surface conditions output by MESA (i.e. , log(), and metallicity ). Furthermore, colors can then convolve these synthetic SEDs with filter transmission curves from astronomical surveys including and , the latter of which we use in this work.
@@
We use the ATLAS9 atmospheric model grid (F. Castelli & R. L. Kurucz, 2003), the data and interpolation for which is provided by default in the colors module. However, it should be noted that this grid does not span the full parameter space to cover that of quasi-stars at all points in their lifetimes. Fig. 3 shows a Kiel diagram containing points for various quasi-star models. The small gray region shows the log() – space covered by the ATLAS9 grid. Although all model temperatures lie firmly within the grid range, many of the quasi-star models lie outside of the log() range. Furthermore, the lowest metallicity provided is . This leaves only the M⊙, quasi-star evolutionary model that fits within the atmosphere grid. However, comparison of the low and zero metallicity models shows very little difference in the final SEDs, and therefore we use a M⊙, model for all SEDs in this work.
@@
A few caveats arise due to our adoption of the ATLAS9 atmosphere models. First, these models are computed in local thermodynamic equilibrium (LTE) and therefore lack the physics necessary to produce all of the emission lines required for rigorous analysis. Additionally, a solar-scaled abundance mixture is assumed, and thus absorption features that would not otherwise be present in metal-free atmospheres are seen. This primarily appears in our models at the H and K singly-ionized calcium (CaII) lines at 3968 Å and 3934 Å respectively. While these strong absorption features are present in the adopted ATLAS9 atmosphere models, they can safely be ignored in this work as there is little to no Ca present in the atmospheres of our quasi-star models, and are expected to be far less abundant in the composition of the gas at the high redshifts studied in this work. It is however conceivable that quasi-stars may have some enrichment via accretion onto the envelope from a host galaxy that has undergone some star formation. We emphasize again that our comparisons are between the continuum features of the models and observations and serve as a pilot study. The modeling and implementation of more detailed atmosphere models will be the subject of future study.
2.3 Host Galaxy & Dust Corrections

### v2

@@ Abstract
JWST has revealed a population of red, compact, high-redshift () objects referred to as “Little Red Dots” (LRDs). These objects exhibit unusual spectral features reminiscent of stellar spectra with blackbody-like SEDs, large hydrogen Balmer breaks, Balmer line absorption, and classical stellar absorption features such as calcium H&K and the calcium triplet. Following the recent suggestion that these may be actively accreting direct-collapse black holes in the process of assembly, i.e. quasi-stars, we present evolutionary models of quasi-stars using our recently released, publicly available MESA-QUEST modeling framework. We compute a grid of models spanning a range of black hole masses and predict the luminosities, temperatures, surface gravities, and lifetimes of these objects. We find that these models lie along a Hayashi track once they hit their “late-stage” which constitutes the majority of their lives ( Myr). We present scaling relations for estimating the mass of a quasi-star as a function of the bolometric luminosity, as well as the bolometric luminosity as a function of the effective temperature for the Hayashi track. The short lifetimes in tandem with the observed number density of LRDs imply the possibility that every supermassive black hole detected at these epochs was once a quasi-star. We compare synthetic spectra of our quasi-star models to observations of LRDs, and show that these models are broadly capable of reproducing the continuum spectra of observed LRDs. These results indicate that quasi-stars are promising candidates for the origin of supermassive black holes via direct collapse in the early universe.
@@ I Introduction
The rapid emergence of supermassive black holes (SMBHs) within the first Myr of cosmic history remains a major puzzle. Observations indicating black hole-to-stellar mass ratios up to , significantly higher than those observed locally (Harikane et al., 2023; Jones et al., 2025; Maiolino et al., 2025; Yue et al., 2024), suggest that the theoretically proposed “direct collapse” pathway for forming massive initial black hole seeds may have been in operation in the early Universe (Whalen et al., 2023; Natarajan et al., 2024; Coughlin and Begelman, 2024). In this scenario, a heavy seed black hole of order M⊙ forms from the gravitational collapse of low metallicity gas in early-universe dense pre-galactic gas clouds in various cosmic settings, which then accretes at the Eddington limit until it fully forms a SMBH (Haehnelt and Rees, 1993; Loeb and Rasio, 1994; Eisenstein and Loeb, 1995; Madau and Rees, 2001; Bromm and Loeb, 2003; Lodato and Natarajan, 2006; Begelman et al., 2006; Natarajan et al., 2017) or from the rapid amplification of a light initial seed that can grow at super-Eddington rates (Devecchi and Volonteri, 2009; Mayer et al., 2010; Alexander and Natarajan, 2014).
@@
The evolution of direct collapse heavy seed formation is expected to proceed in the form of a quasi-star (Lodato and Natarajan, 2006; Begelman et al., 2008). These are formed when the core of either a supermassive star or a pre-galactic gas cloud directly collapses initially into a stellar mass black hole around which the remaining gas is assembled as an envelope. This envelope around the growing black hole seed is then supported by the radiation pressure produced from accretion onto the black hole, which can rapidly accrete at super-Eddington rates due to the efficient transport of energy via convection and jets, allowing the central black hole to accrete at roughly the Eddington limit of the entire quasi-star (Begelman et al., 2006, 2008; Begelman, 2010; Ball et al., 2012; Fiacconi and Rossi, 2015; Campbell et al., 2025).
@@
Among the high-redshift, actively-accreting black hole observations by JWST are a subset of candidate active galactic nuclei (AGN) that have come to be known as “Little Red Dots” (LRDs) – high-redshift, morphologically compact, ultra-luminous objects with unique spectral characteristics (Matthee et al., 2024; Greene et al., 2024; labbe25; Akins et al., 2024; Juodžbalis et al., 2024; Kokorev et al., 2024; Kocevski et al., 2025; Hviding, Raphael E. et al., 2025; Ji et al., 2025; Taylor et al., 2025). These objects exhibit an unusually strong Balmer break as well as an extended, flat spectrum in the near-infrared (NIR). The combination of these features, in addition to a blue UV slope, results in a unique “V”-shaped spectral energy distribution (SED, Barro et al., 2024; Killi, Meghana et al., 2024; de Graaff, Anna et al., 2025; Kocevski et al., 2025; Labbe et al., 2023).
@@
Currently, there is no complete explanation for what these LRDs are composed of and hence their unique signatures. The relative contributions of emitted flux originating from star formation versus accretion onto a central BH are debated. Nandal and Loeb (2025) suggest that they can be explained by supermassive stars due to their high luminosities and low photospheric temperatures. Zwick et al. (2025) add on to this and suggest the supermassive star would be embedded within a self-gravitating accretion disc formed from gas-rich galaxy mergers in order to explain the two-blackbody, V-shaped continuum. Another explanation put forth by several studies suggests that LRDs are black holes embedded in dense gas, with Naidu et al. (2025) dubbing their model a “black hole star” (e.g., de Graaff, Anna et al., 2025; Inayoshi and Maiolino, 2025; Ji et al., 2025; Kido et al., 2025; Liu et al., 2025; Taylor et al., 2025). This is motivated by their nearly blackbody-like SEDs, prominent hydrogen Balmer breaks, Balmer line absorption, and classical stellar absorption features such as Ca H&K and the Ca II triplet, which are consistent with emission from an optically thick, thermally emitting envelope surrounding an accreting black hole. Jeon et al. (2025) further proposed that LRDs could be remnants of direct collapse, consistent with the quasi-star scenario (Begelman et al., 2006, 2008; Begelman, 2010), and Begelman and Dexter (2025) have argued that they would most likely represent the “late stage” of this process where the BH has accreted at least 10% the total mass of the system.
@@
Thus far, there has been no comparison between quasi-star evolutionary models and JWST’s LRDs. In this work, we present simulations of quasi-star evolution using MESA (Paxton2011; Paxton2013; Paxton2015; Paxton2018; Paxton2019; Jermyn2023) and the MESA-QUEST modeling framework (Campbell et al., 2025; Santarelli et al., 2025). We extract spectral predictions to compare with recent JWST observations of three LRDs: “The Cliff” (de Graaff, Anna et al., 2025), MoM-BH*-1 (Naidu et al., 2025), and UNCOVER-45924 (Labbe et al., 2024). We compare our predictions with observations for both a quasi-star by itself and a quasi-star embedded within a star-forming galaxy, akin to the scenario predicted by Natarajan et al. (2017). We also consider simple corrections for dust attenuation, but find little evidence for it, in line with recent findings that dust attenuation in LRDs may be subtle or negligible (Casey et al., 2025; de Graaff, Anna et al., 2025; Naidu et al., 2025; Setton et al., 2025; Xiao, Mengyuan et al., 2025). We find that our fiducial M⊙ quasi-star model broadly reproduces the continuum spectra of these LRDs, thereby lending evidence that SMBHs may assemble by passing through the direct-collapse quasi-star phase.
@@
The outline of our paper is as follows: in Section II, we provide details of how we extract spectra from quasi-star models in MESA-QUEST. We present the results from models and comparisons with observational data of three JWST LRDs in Section III, and close with a final discussion and implications of these results.
@@
II Methods: extracting quasi-star spectra from MESA-QUEST
In this section, we briefly describe our quasi-star MESA models.
@@ II Methods: extracting quasi-star spectra from MESA-QUEST
We use our previously developed MESA-QUEST framework, which includes several newly implemented upgrades: flexible inner boundary conditions, accretion schemes, and improvements that enable modeling higher total masses than previously studied.
The following section briefly outlines these methods; more details can be found in Santarelli et al. (2025) and Campbell et al. (2025).
We note that all MESA work directories are publicly available on Zenodo with a living version available via the MESA-QUEST Github repository.111Zenodo: https://doi.org/10.5281/zenodo.17992486 (catalog 10.5281/zenodo.17992486)
GitHub: https://github.com/andysantarelli/MESA-QUEST
II.1 Quasi-star Modeling
@@
A quasi-star is modeled as a black hole enveloped within a star via an adjustment to the inner boundary conditions of the stellar structure equations such that:
@@
where is the innermost grid point, is the updated inner boundary, and and are the black hole mass and accretion luminosity respectively. In this work, we use the “saturated-convection” radius, the boundary within which material is slowly drifting inward while convection transports energy outward at the highest possible rate (Coughlin and Begelman, 2024). Details on the calculation of this boundary can be found in Coughlin and Begelman (2024), with further implementation details in Santarelli et al. (2025) and Campbell et al. (2025). For the accretion luminosity, and thus the accretion rate, we use the Eddington limit of the entire object such that
@@
Previous works modeling the accretion and resulting luminosity report only a factor of deviation from the object’s Eddington limit throughout the quasi-star’s lifetime (Ball et al., 2012; Campbell et al., 2025). The dimensionless scaling factor serves as a proxy for a range of physical effects that enhance or suppress accretion such as rotation, magnetic fields, photon-trapping, etc., which we plan to implement and explore more thoroughly in future work. Here we adopt a simple scaling factor and use as a first approximation.
@@
after which the quasi-star “settles” as the black hole mass reaches M⊙. It is unclear whether this initial phase is physical or a result of the initialization of the model. The inset plot shows the majority of the quasi-star’s life.
The final age of this model is Myr, i.e., the timescale for which a 10 M⊙ black hole accretes at the Eddington limit of the entire object until . The actual lifetimes may be longer as our models do not currently indicate what happens once the inner boundary condition exceeds the total radius, and it has been suggested that the envelope would be ejected once the surface temperature cools to K (Begelman et al., 2008; Ball et al., 2012). It is unclear when or if this would occur after our model is terminated. If we assume that BH growth continues at the Eddington rate of the entire object until the envelope is completely consumed, we expect a total lifetime of Myr.
@@
where for convenience we have dropped the dependency on and , though we note that their variation can have a large influence on the accretion physics and thus the luminosity itself (Santarelli et al., 2025).
Within mixing length theory, the mixing length parameter is a dimensionless parameter that defines the distance that a parcel of convective material can travel in terms of the pressure scale height (Böhm-Vitense, 1958; Cox and Giuli, 1968). While MESA uses a default value of , is a free parameter of order unity that should vary across models. We show in the left plot of Fig. 4 the impact that it has on our models: decreasing results in an increase in radius and a lower , and thus a reddening of the SED. For this work, we choose and for our M⊙ total mass models as they best fit the SEDs of the selected LRDs. While models with masses of M⊙ have similar effective temperatures with higher values of , they do not have the bolometric luminosities seen in most LRDs. Further details on extraction of the SEDs from the models as well as the LRD comparisons can be found in Sections II.2 and III.
A detailed analysis of the impact of our modeling assumptions (e.g. accretion rate, boundary conditions, surface winds) can be found in Santarelli et al. (2025) with additional comparison of a similar model to previous analytic models in Hassan et al. (in prep.-10-06). Santarelli et al. (2025) shows that changes to the accretion rate, particularly for accretion stifling cases, may result in smaller radii and thus hotter surface temperatures. The inclusion of surface winds is shown to have a large impact on the stability of the quasi-star and thus the final black hole mass. Both the black hole accretion physics and the offset to the wind-driven mass loss from envelope accretion remain uncertain and will be studied in future work.
We have explored the inclusion of sparse metals within the envelope as well as alternate opacity tables. Changes to either of these parameters results in no qualitative difference in the evolution or emergent spectra discussed in the following section. The alternate opacity tables have negligible impact due to the efficient convection in the envelope. For models with increased metallicity, the similarities in spectra are likely due to the limited parameter space covered by the adopted atmospheric grids (see Sec. II.2 and Fig. 3).
II.2 Extracting Synthetic Spectral-Energy Distributions
@@
In order to extract the synthetic spectra of quasi-stars, we use the MESA colors module, developed by Miller (2025). This is a built-in tool within MESA that can be used to generate synthetic spectral-energy distributions (SEDs) by sampling the nearest grid point in a stellar atmosphere model grid using the stellar surface conditions output by MESA (i.e. , log(), and metallicity ). Furthermore, colors can then convolve these synthetic SEDs with filter transmission curves from astronomical surveys including and , the latter of which we use in this work.
@@
We use the ATLAS9 atmospheric model grid (Castelli and Kurucz, 2003), the data and interpolation for which is provided by default in the colors module. However, it should be noted that this grid does not span the full parameter space to cover that of quasi-stars at all points in their lifetimes. Fig. 3 shows a Kiel diagram containing points for various quasi-star models, all with . The small gray region shows the log() – space covered by the ATLAS9 grid. Although all model temperatures lie firmly within the grid range, the models lie outside of the log() range. Furthermore, the lowest metallicity provided is . However, comparison of the low and zero metallicity models shows very little difference in the final SEDs and log() is off of the grid by in log-space for the M⊙ model with , and when . Therefore we use a M⊙, model for all SEDs in this work.
@@
A few caveats arise due to our adoption of the ATLAS9 atmosphere models. First, these models are computed in local thermodynamic equilibrium (LTE) and therefore lack the physics necessary to produce all of the emission lines required for rigorous analysis. Additionally, a solar-scaled abundance mixture is assumed, and thus absorption features that would not otherwise be present in metal-free atmospheres are seen. This appears primarily in our models on the H and K singly-ionized calcium (CaII) lines at 3968 Å and 3934 Å respectively. While these strong absorption features are present in the adopted ATLAS9 atmosphere models, they can safely be ignored in this work, as there is little to no Ca present in the atmospheres of our quasi-star models and are expected to be far less abundant in the composition of the gas at the high redshifts studied in this work than in the solar-scaled abundances in the atmosphere models. It is, however, conceivable that quasi-stars may have some enrichment via accretion onto the envelope from a host galaxy that has undergone some star formation. We emphasize again that our comparisons are between the continuum features of the models and observations and serve as a pilot study. The modeling and implementation of more detailed atmosphere models will be the subject of future study.
II.3 Host Galaxy & Dust Corrections